OpenStudy (anonymous):

If what makes an acid strong is the extent of dissociation, at what sorts of concentrations (high or low) do weak acids behave more like strong acids? Explain

OpenStudy (anonymous):

Think of your equilibrium problem:\[\frac{[{\rm H}^+][{\rm A}^-]}{[{\rm HA}]} = K_a\]As long as [H+] >> 10^-7, [H+] = [A-] (all the H+ will come from the dissociation of the weak acid), and %dissoc = [A-]/[HA] * 100, so you can rewrite this as \[[{\rm HA}] \times \left( 100 \times {\rm \%dissoc} \right)^2 = K_a\]Or\[({\rm \%dissoc})^2 = \frac{K_a}{10^4} \frac{1}{[{\rm HA}]}\]In short, the % dissociation is inversely proporational to the concentratio of the weak acid. Lower concentrations mean higher dissociation -- more like a strong acid.